OpenStudy (anonymous):

3618.85 = 3.13x + 340 sqrt(2x) Solve for x

OpenStudy (anonymous):

This might seem tricky at first but the trick to the question is simply "changing" the sqrt(2x) into something we are more familiar with. Let's break down this question. Let sqrt(2x) = u (2x) = u^2 (x) = (u^2)/ 2 Now that we've simplified it, just simply plug this all back into the original question. 3618.85 = 3.13 (u^2)/ 2 + 340 u 3618.85 = 1.565 u^2 + 340 u Solve the simple quadratic by using a handy dandy calculator and your answer for u should be u = 10.1678 and some other negative number, however we omit the negative one because the square root of a negative number does not exist. Sub the u value in to get the original x value and x should equal 51.6921.

OpenStudy (anonymous):

Hope this was easy to follow :3